# Macroscopic properties of equity markets: stylized facts and portfolio performance

### v1

@@ 4.1. Excess growth rate
    The excess growth rate is numéraire invariant, i.e., $\gamma_{w}$ remains unchanged if the asset prices, and hence their returns, are measured with respect to another numériare such as a benchmark portfolio [91, Lemma 3.2]. This makes the excess growth rate an appropriate measure of relative volatility. [^13] Numéraire invariance uniquely characterizes the excess growth rate among a class of distance-like quantities called $L$-divergences; see [93, Example 3.10].
@@
    which is the discrete analogue of (A.1). The left hand side of (4.4) is the log-return $r_{w}(T_{0},T_{1})$ of the portfolio rebalanced to the constant weight $w$ at each time $t_{\ell}$ (here and in (iii) we neglect transaction costs). Everything else equal, a large relative volatility is favourable to a rebalanced portfolio compared to a capitalization-weighted one. This observation underlies the idea of volatility pumping [16, 37, 91, 97] and can be formalized mathematically using the concept of functional portfolio generation in SPT [46, 49, 92, 114]. [^14] For a general portfolio (not constant weighted or functionally generated), pathwise decompositions such as (4.4) do not apply and a direct attribution of performance to its excess growth rate is less clear. See [24, 111] and the references therein for discussions about the limitations of the excess growth rate or diversifcation return in explaining “alpha”. In this section, we simply use the excess growth rate as a measure of intrinsic volatility. In Section 6 we relate the excess growth rate with the performance of the diversity-weighted portfolio which is functionally generated. Thus it is natural to consider maximization of excess growth rate subject to suitable constraints [46, Example 1.1.7]. For example, risk-adjusted performance of portfolios optimized with respect to excess growth rate was investigated in [86], and a generalized efficient frontier was constructed in [38].
@@
    where $\tilde{\Gamma}_{w}$ is the cumulated excess growth rate with respect to the time grid $\{\tilde{t}_{m}\}$. Thus the spread in (4.5) captures the effect of the frequency of rebalancing. The authors of [56] observed in a high-frequency setting that the cumulative excess growth rate increases as the frequency of rebalancing increases and devised a long-short portfolio to exploit the spread.
@@ 5. Rank-based properties
In this section we study how the behaviours of stocks depend systematically on their relative ranks with respect to market capitalization. [^18] This is analogous to the practice of forming deciles in empirical finance: for each period, group the assets after ordering them with respect to some criterion. Here the criterion is rank (by market capitalization) and we will examine quantities indexed by rank. In stochastic portfolio theory there are two main motivations for the study of rank-based properties. First, as mentioned in Section 3, “projecting” to the space of ranked market capitalization or weights (thus neglecting name-based features as in the rank-based diffusion system (3.3)) allows for more parsimonious modelling of some features of market capitalizations and the capital distribution curve. Second, it is common to restrict the investment universe to a top segment of stocks by market capitalization (one reason is that these stocks are more representative of the market and are also more liquid). When strictly enforced, this means selling a stock when it drops out of the universe and buying a stock when it enters. Thus part of the turnover of the portfolio is directly related to the “intensity” of rank switching near the cut off rank. [^19] Rank switching also plays an important role when the benchmark is a capitalization-weighted portfolio consisting of the largest $K$ stocks (rather than the entire market). Leakage refers to the loss in wealth due to renewing the portfolio’s constituent stocks. We refer the reader to [46, Example 4.3.5] and [115] for a detailed discussion of leakage and its computation.
@@ 5.3. Intensity of rank switching
In Figure 12 we plot the intensity $\Lambda_{k}(t)$ as a function of time, for $k=1,\ldots,499$. We observe $\Lambda_{k}(t)$ increases roughly linearly in time and in $k$. Moreover, the intensity of rank switching is very small for the largest stocks ($k$ small). As seen in Figure 6, the largest stocks mostly remain on top and rank switches only occur occasionally. These findings are consistent with the results (concerning the period 19901990199019901999199919991999) presented in [46, Section 5.2].
@@ 6.1. Set-up of the experiment
    In (ii) and (iii), all trading activities (except initialization) incur proportional transaction costs following exactly the procedure described in [100, Section 2.1]. In view of [18, 100], we choose a cost of $0.25\%$ for all stocks on purchases and sales. The empirical study [5] and the NASDAQ report [84] also support a choice of cost on this order of magnitude.
@@ 6.2.2. Sensitivity to portfolio specifications
The top right panel of Figure 14 plots the annual performance of the equal weight portfolio $w^{0,f}$ relative to the buy-and-hold (initially equal-weighted) portfolio $w^{0,\infty}$ for various fixed rebalancing frequencies, $f$. Since there are approximately 250 trading days every year, we choose the set of divisors $\{1,2,5,10,25,50,125,\infty\}$ [^27] If a year has 250 trading days then $f=\infty$ will give the same result as any $f>250$.. The range of the gray bars represents the relative performance spanned by the portfolios in this set. Once again, by construction, when $f=\infty$ the relative performance is 00. Within each of the bars we mark the value corresponding to a given $f$ by a colored line that changes on gradient from blue to red as $f$ increases. Unlike the top left panel, the relationship here is not monotone in general. There are many years when intermediate values of $f$ outperform. We see that, for most years, the choice of frequency is less material than the choice of $p$ in determining the degree of over/under-performance. That said, we observe that frequent rebalancing (say, every 1–2 days) underperforms in most years under $0.25\%$ transaction cost. Also, around the financial crises of 2000 and 2008 the right rebalancing decision appears to meaningfully impact portfolio returns. The bottom right panel displays the difference in the cumulative daily excess growth rate and the annual excess growth rate for the calendar years in our backtest. This statistic captures the difference in the intrinsic volatility at these two different rebalancing scales. However, we see that this only has a weak positive relationship with the difference in performance due to changes in $f$. This is not unexpected, since the contribution of the excess growth rate to portfolio returns is a second order effect (see, e.g. Figure 13) and the strength of the relationship can be confounded by many other market factors including transaction costs, defaults, and a variable set of constituent securities.
@@ 6.2.3. Sharpe ratio and drawdown
To date, relatively little has been studied about portfolio risks and risk adjusted returns in the context of stochastic portfolio theory. In their empirical analysis, [100] reported the historical Sharpe ratios of several SPT portfolios and illustrated an approximately linear decay in increasing transaction costs. There has also been some related theoretical work on drawdown constrained growth rate optimization (see [35] and [73, Section 4.1]). The first and last author proposed in [19] an optimization framework that allows flexibility in the exposure to market diversity. In view of this gap in the literature, we take the opportunity to investigate the risk profile of the diversity-weighted portfolios $w^{p,f}$ that have been the subject of our empirical investigation.
@@ 7. Conclusion
    (Other quantities and intraday data) In this paper we focused on macroscopic properties defined in terms of market capitalizations and returns. Other quantities, such as trading volumes, signatures (in the sense of rough path theory) [31], sectors, and prices of options and exchange traded funds, may offer additional insights. Also, the CRSP Database offers daily but not intraday data. SPT-inspired analysis of high frequency data is promising but largely open; the only work we are aware of is [56] which studied intraday properties of excess growth rate in relation to the profits made by market makers. A possible direction is to study intraday return distributions (see e.g. [116]) and relate them with market capitalizations and other quantities.
@@ Acknowledgement
S. Campbell and T.-K. L. Wong thank Martin Larsson, David Itkin and Robert Jones for helpful discussions. The research of T.-K. L. Wong is partially supported by an NSERC Discovery Grant (RGPIN-2019-04419) and a Seed Funding for Methodologists Grant from the Data Sciences Institute (DSI) at the University of Toronto.
@@ Appendix A Crash course in stochastic portfolio theory
Consider an equity market model consisting of a fixed collection of $n$ non-dividend paying stocks. [^29] Dividends are included in Fernholz’s original market model [46] but are frequently omitted for convenience. See [12, 74] for recent attempts to relax the assumption of a fixed investment universe. The market capitalizations $X_{1}(t),\ldots,X_{n}(t)$ of the stocks are modelled as a vector of positive Itô processes defined on some filtered probability space. The vector $\mu(t)=(\mu_{1}(t),\ldots,\mu_{n}(t))$ of market weights, where $\mu_{i}(t)=X_{i}(t)/(X_{1}(t)+\cdots+X_{n}(t))$, takes values in the open unit simplex $\Delta_{n}$. The diversity of the market refers to the concentration of the capital distribution, and may be quantified in terms of a symmetric concave function $\Phi:\Delta_{n}\rightarrow(0,\infty)$ (see [46, Section 3.4]). Popular choices include the Shannon entropy ${\bf H}(\mu)$ (see (3.4)) as well as the parameterized diversity ${\bf D}_{p}(\mu)$ (see (3.5)). It was observed empirically that market diversity tends to “mean-reverting” in the long run (see Section 3.2). At the very least, the market weight vector $\mu(t)$ tends to avoid certain regions of the simplex especially its vertices. Motivated by this and other empirical observations, various path properties may be imposed on the market model. For example, the market is said to be coherent if a.s. $\lim_{t\rightarrow\infty}\frac{1}{t}\log\mu_{i}(t)=0$ for all $i$, and is diverse if there exists $\delta>0$ such that a.s. $\max_{1\leq i\leq n}\mu_{i}(t)\leq 1-\delta$ for all $i$ and $t\geq 0$. Note that some conditions, when imposed to hold with probability $1$, exclude the existence of an equivalent martingale measure (over a finite horizon); see [55, Section 6] for a discussion.
@@
If the market is coherent, the first term on the right hand side of (A.6) tends to 00 as $t\rightarrow\infty$. Suppose further that the market is diverse and is nondegenerate in the sense that the covariance process of $\log X(t)$ is uniformly elliptic [46, Definition 1.1.2]. Then, it can be shown that $\Gamma_{\pi}(t)\geq\epsilon t$ for some constant $\epsilon>0$. Letting $t\rightarrow\infty$ in (A.6), we have $\liminf_{t\rightarrow\infty}\frac{1}{t}\log\frac{V_{\pi}(t)}{V_{\pi}(0)}\geq\epsilon$ a.s., thus giving an “asymptotic relative arbitrage”.
To give an explicit example of relative arbitrage in the sense of (A.5), consider the diversity-weighted portfolio which is closely related to the function (3.5) and is given by
@@
From (A.7), the relative performance of the portfolio is characterized by (i) the change in market diversity and (ii) market volatility measured by the excess growth rate of the portfolio. The excess growth rate contributes positively to relative performance, but over the short run the dynamics of $\log V_{\pi}(t)$ is dominated by that of $\log{\bf D}_{p}(\mu(t))$. As the parameter $p$ varies between 00 and $1$, we have a trade-off between exposure to intrinsic volatility versus diversity. [^30] See [109] for an analysis of the diversity-weighted portfolio when $p$ is negative. Assume that the market is diverse, so that $\log{\bf D}_{p}(\mu(t))\geq-M$ for some $M>0$. If the market is also nondegenerate, then $(1-p)\Gamma_{\pi}(t)\geq\epsilon t$ for some $\epsilon>0$. This gives the lower bound
@@
and we have a relative arbitrage over $[0,T]$ whenever $T>T^{*}:=\frac{1}{\epsilon}(M+\log{\bf D}_{p}(\mu(0)))$ (unfortunately, $T^{*}$ may be too large for practical purposes). Both the constant-weighted portfolio and the diversity-weighted portfolio are special cases of functionally generated portfolios for which a pathwise decomposition analogous to (A.7) holds. [^31] Analogous pathwise decompositions can be established without a stochastic model. See [92] for the discrete time case and [3] for the continuous time case using rough path theory. For generalizations of functional portfolio generalization see [75, 78, 114] and the references therein. Relative arbitrages can also be constructed over short time horizons or under weaker conditions. For further details and more recent developments, we refer the reader to [3, 19, 47, 74, 92] and the references therein.
@@ Appendix B Local time
We recall the concept of local time in stochastic calculus which motivates the development in Section 5.3 and is closely related to the concept of leakage in stochastic portfolio theory. We follow the notations of [46, Section 4.1]. For further details see [98, Chapter VI], [9] and [115]. Let $Y(t)$ be a real-valued continuous semimartingale defined on a filtered probability space. The local time of $Y$ at 00 is the non-decreasing process $\Lambda$ defined by the identity
@@
[46, Proposition 4.1.11] Let $Y_{1},\dots,Y_{n}$ be pathwise mutually non-degenerate and absolutely continuous. Then $(Y_{(1)},\ldots,Y_{(n)})$ is an $n$-dimensional continuous semimartingale. For $k=1,\ldots,n-1$, let $\Lambda_{k}(t)$ be the local time of the non-negative semimartingale $Y_{(k)}-Y_{(k+1)}$ at 00. Also let $\Lambda_{0}(t)=\Lambda_{n+1}(t)=0$. Then

### v2

@@ 4.1. Excess growth rate
    The excess growth rate is numéraire invariant, i.e., $\gamma_{w}$ remains unchanged if the asset prices, and hence their returns, are measured with respect to another numériare such as a benchmark portfolio [96, Lemma 3.2]. This makes the excess growth rate an appropriate measure of relative volatility. [^14] Numéraire invariance uniquely characterizes the excess growth rate among a class of distance-like quantities called $L$-divergences; see [98, Example 3.10].
@@
    which is the discrete analogue of (A.1). The left hand side of (4.4) is the log-return $r_{w}(T_{0},T_{1})$ of the portfolio rebalanced to the constant weight $w$ at each time $t_{\ell}$ (here and in (iii) we neglect transaction costs). Everything else equal, a large relative volatility is favourable to a rebalanced portfolio compared to a capitalization-weighted one. This observation underlies the idea of volatility pumping [16, 38, 96, 102] and can be formalized mathematically using the concept of functional portfolio generation in SPT [47, 50, 97, 119]. [^15] For a general portfolio (not constant weighted or functionally generated), pathwise decompositions such as (4.4) do not apply and a direct attribution of performance to its excess growth rate is less clear. See [25, 116] and the references therein for discussions about the limitations of the excess growth rate or diversifcation return in explaining “alpha”. In this section, we simply use the excess growth rate as a measure of intrinsic volatility. In Section 6 we relate the excess growth rate with the performance of the diversity-weighted portfolio which is functionally generated. Thus it is natural to consider maximization of excess growth rate subject to suitable constraints [47, Example 1.1.7]. For example, risk-adjusted performance of portfolios optimized with respect to excess growth rate was investigated in [91], and a generalized efficient frontier was constructed in [39].
@@
    where $\tilde{\Gamma}_{w}$ is the cumulated excess growth rate with respect to the time grid $\{\tilde{t}_{m}\}$. Thus the spread in (4.5) captures the effect of the frequency of rebalancing in the absence of transaction costs. The authors of [57] observed in a high-frequency setting that the cumulative excess growth rate increases as the frequency of rebalancing increases and devised a long-short hedging portfolio to exploit the spread.
@@ 5. Rank-based properties
In this section we study how the behaviours of stocks depend systematically on their relative ranks with respect to market capitalization. [^19] This is analogous to the practice of forming deciles in empirical finance: for each period, group the assets after ordering them with respect to some criterion. Here the criterion is rank (by market capitalization) and we will examine quantities indexed by rank. In stochastic portfolio theory there are two main motivations for the study of rank-based properties. First, as mentioned in Section 3, “projecting” to the space of ranked market capitalization or weights (thus neglecting name-based features as in the rank-based diffusion system (3.3)) allows for more parsimonious modelling of some features of market capitalizations and the capital distribution curve. Second, it is common to restrict the investment universe to a top segment of stocks by market capitalization (one reason is that these stocks are more representative of the market and are also more liquid). When strictly enforced, this means selling a stock when it drops out of the universe and buying a stock when it enters. Thus part of the turnover of the portfolio is directly related to the “intensity” of rank switching near the cut off rank. [^20] Rank switching also plays an important role when the benchmark is a capitalization-weighted portfolio consisting of the largest $K$ stocks (rather than the entire market). Leakage refers to the loss in wealth due to renewing the portfolio’s constituent stocks. We refer the reader to [47, Example 4.3.5] and [120] for a detailed discussion of leakage and its computation.
@@ 5.3. Intensity of rank switching
In Figure 12 we plot the intensity $\Lambda_{k}(t)$ as a function of time, for $k=1,\ldots,499$. We observe $\Lambda_{k}(t)$ increases roughly linearly in time and in $k$. Moreover, the intensity of rank switching is very small for the largest stocks ($k$ small). As seen in Figure 6, the largest stocks mostly remain on top and rank switches only occur occasionally. These findings are consistent with the results (concerning the period 19901990199019901999199919991999) presented in [47, Section 5.2].
@@ 6.1. Set-up of the experiment
    In (ii) and (iii), all trading activities (except initialization) incur proportional transaction costs following exactly the procedure described in [105, Section 2.1]. In view of [18, 105], we choose a cost of $0.25\%$ for all stocks on purchases and sales. The empirical study [5] and the NASDAQ report [89] also support a choice of cost on this order of magnitude.
@@ 6.2.2. Sensitivity to portfolio specifications
The top right panel of Figure 14 plots the annual performance of the equal weight portfolio $w^{0,f}$ relative to the buy-and-hold (initially equal-weighted) portfolio $w^{0,\infty}$ for various fixed rebalancing frequencies, $f$. Since there are approximately 250 trading days every year, we choose the set of divisors $\{1,2,5,10,25,50,125,\infty\}$ [^28] If a year has 250 trading days then $f=\infty$ will give the same result as any $f>250$.. The range of the gray bars represents the relative performance spanned by the portfolios in this set. Once again, by construction, when $f=\infty$ the relative performance is 00. Within each of the bars we mark the value corresponding to a given $f$ by a colored line that changes on a gradient from blue to red as $f$ increases. Unlike the top left panel, the relationship here is not monotone in general. There are many years when intermediate values of $f$ outperform. We see that, for most years, the choice of frequency is less material than the choice of $p$ in determining the degree of over/under-performance. That said, we observe that frequent rebalancing (say, every 1–2 days) underperforms in most years under $0.25\%$ transaction cost. Also, around the financial crises of 2000 and 2008 the right rebalancing decision appears to meaningfully impact portfolio returns. The bottom right panel displays the difference in the cumulative daily excess growth rate and the annual excess growth rate for the calendar years in our backtest. This statistic captures the difference in the intrinsic volatility at these two different rebalancing scales. However, we see that this only has a weak positive relationship with the difference in performance due to changes in $f$. This is not unexpected, since the contribution of the excess growth rate to portfolio returns is a second order effect (see, e.g. Figure 13) and the strength of the relationship can be confounded by many other market factors including transaction costs, defaults, and a variable set of constituent securities.
@@ 6.2.3. Sharpe ratio and drawdown
To date, relatively little has been studied about portfolio risks and risk adjusted returns in the context of stochastic portfolio theory. In their empirical analysis, [105] reported the historical Sharpe ratios of several SPT portfolios and illustrated an approximately linear decay in increasing transaction costs. There has also been some related theoretical work on drawdown constrained growth rate optimization (see [36] and [77, Section 4.1]). The first and last author proposed in [19] an optimization framework that allows flexibility in the exposure to market diversity. In view of this gap in the literature, we take the opportunity to investigate the risk profile of the diversity-weighted portfolios $w^{p,f}$ that have been the subject of our empirical investigation.
@@ 7. Conclusion
    (Other quantities and intraday data) In this paper we focused on macroscopic properties defined in terms of market capitalizations and returns. Other quantities, such as trading volumes, signatures (in the sense of rough path theory) [32], sectors, prices of options [^30] See [60] for some relationships between contingent claims and functionally generated portfolios in SPT. and exchange traded funds, may offer additional insights. Also, the CRSP Database offers daily but not intraday data. SPT-inspired analysis of high frequency data is promising but largely open; the only work we are aware of is [57] which studied intraday properties of the excess growth rate in relation to the profits made by market makers. A possible direction is to study intraday return distributions (see e.g. [121]) and relate them with market capitalizations and other quantities.
@@ Acknowledgement
S. Campbell and T.-K. L. Wong thank Martin Larsson, David Itkin and Robert Jones for helpful discussions. The research of T.-K. L. Wong is partially supported by an NSERC Discovery Grant (RGPIN-2019-04419) and a Seed Funding for Methodologists Grant from the Data Sciences Institute (DSI) at the University of Toronto. The authors report there are no competing interests to declare.
@@ Appendix A Crash course in stochastic portfolio theory
Consider an equity market model consisting of a fixed collection of $n$ non-dividend paying stocks. [^31] Dividends are included in Fernholz’s original market model [47] but are frequently omitted for convenience. See [12, 78] for recent attempts to relax the assumption of a fixed investment universe. The market capitalizations $X_{1}(t),\ldots,X_{n}(t)$ of the stocks are modelled as a vector of positive Itô processes defined on some filtered probability space. The vector $\mu(t)=(\mu_{1}(t),\ldots,\mu_{n}(t))$ of market weights, where $\mu_{i}(t)=X_{i}(t)/(X_{1}(t)+\cdots+X_{n}(t))$, takes values in the open unit simplex $\Delta_{n}$. The diversity of the market refers to the concentration of the capital distribution, and may be quantified in terms of a symmetric concave function $\Phi:\Delta_{n}\rightarrow(0,\infty)$ (see [47, Section 3.4]). Popular choices include the Shannon entropy ${\bf H}(\mu)$ (see (3.4)) as well as the parameterized diversity ${\bf D}_{p}(\mu)$ (see (3.5)). It was observed empirically that market diversity tends to “mean-reverting” in the long run (see Section 3.2). At the very least, the market weight vector $\mu(t)$ tends to avoid certain regions of the simplex especially its vertices. Motivated by this and other empirical observations, various path properties may be imposed on the market model. For example, the market is said to be coherent if a.s. $\lim_{t\rightarrow\infty}\frac{1}{t}\log\mu_{i}(t)=0$ for all $i$, and is diverse if there exists $\delta>0$ such that a.s. $\max_{1\leq i\leq n}\mu_{i}(t)\leq 1-\delta$ for all $i$ and $t\geq 0$. Note that some conditions, when imposed to hold with probability $1$, exclude the existence of an equivalent martingale measure (over a finite horizon); see [56, Section 6] for a discussion.
@@
If the market is coherent, the first term on the right hand side of (A.6) tends to 00 as $t\rightarrow\infty$. Suppose further that the market is diverse and is nondegenerate in the sense that the covariance process of $\log X(t)$ is uniformly elliptic [47, Definition 1.1.2]. Then, it can be shown that $\mathrm{d}\Gamma_{\pi}(t)\geq\epsilon\mathrm{d}t$ for some constant $\epsilon>0$. Letting $t\rightarrow\infty$ in (A.6), we have $\liminf_{t\rightarrow\infty}\frac{1}{t}\log\frac{V_{\pi}(t)}{V_{\pi}(0)}\geq\epsilon$ a.s., thus giving an “asymptotic relative arbitrage”.
To give an explicit example of relative arbitrage in the sense of (A.5), consider the diversity-weighted portfolio which is closely related to the function ${\bf D}_{p}$ in (3.5) and is given by
@@
From (A.7), the relative performance of the portfolio is characterized by (i) the change in market diversity and (ii) market volatility measured by the excess growth rate of the portfolio. The excess growth rate contributes positively to relative performance, but over the short run the dynamics of $\log V_{\pi}(t)$ is dominated by that of $\log{\bf D}_{p}(\mu(t))$. As the parameter $p$ varies between 00 and $1$, we have a trade-off between exposure to intrinsic volatility versus diversity. [^32] See [114] for an analysis of the diversity-weighted portfolio when $p$ is negative. Assume that the market is diverse, so that $\log{\bf D}_{p}(\mu(t))\geq-M$ for some $M>0$. If the market is also nondegenerate, then $\mathrm{d}(1-p)\Gamma_{\pi}(t)\geq\epsilon\mathrm{d}t$ for some $\epsilon>0$. This gives the a.s. lower bound
@@
and we have a relative arbitrage over $[0,T]$ whenever $T>T^{*}:=\frac{1}{\epsilon}(M+\log{\bf D}_{p}(\mu(0)))$ (unfortunately, the bound $T^{*}$ obtained this way is too large for practical purposes). Both the constant-weighted portfolio and the diversity-weighted portfolio are special cases of functionally generated portfolios for which a pathwise decomposition analogous to (A.7) holds. [^33] Analogous pathwise decompositions can be established without a stochastic model. See [97] for the discrete time case and [3] for the continuous time case using rough path theory. For generalizations of functional portfolio generalization see [79, 83, 119] and the references therein. Relative arbitrages can also be constructed over short time horizons or under weaker conditions. For further details and more recent developments, we refer the reader to [3, 19, 48, 78, 97] and the references therein.
@@ Appendix B Local time
We recall the concept of local time in stochastic calculus which motivates the development in Section 5.3 and is closely related to the concept of leakage in stochastic portfolio theory. We follow the notations of [47, Section 4.1]. Further details can be found in [103, Chapter VI], [9] and [120]. Let $Y(t)$ be a real-valued continuous semimartingale defined on a filtered probability space. The local time of $Y$ at 00 is the non-decreasing process $\Lambda$ defined by the identity
@@
[47, Proposition 4.1.11] Let $Y_{1},\dots,Y_{n}$ be pathwise mutually non-degenerate and absolutely continuous. Then $(Y_{(1)},\ldots,Y_{(n)})$ is an $n$-dimensional continuous semimartingale. For $k=1,\ldots,n-1$, let $\Lambda_{k}(t)$ be the local time of the non-negative semimartingale $Y_{(k)}-Y_{(k+1)}$ at 00. Also let $\Lambda_{0}(t)=\Lambda_{n+1}(t)=0$. Then